solve cos(x/2)=1/2, 0 less than or equal to x less than 2 pi
Answer by lwsshak3(11628)

solve cos(x/2)=1/2, 0 less than or equal to x less than 2 pi
cos(x/2)=1/2
x/2=π/3,5π/3 (in quadrants I and IV where cos>0
x=2π/3,10π/3